# Average Certified Radius is a Poor Metric for Randomized Smoothing

### v2

@@ 4.3 Selection Bias in Existing RS Training Algorithms
We have shown that ACR strongly prefers easy samples in Section 4.2. However, since ACR is not differentiable with respect to the model parameters because it is based on counting, RS training strategies usually do not directly apply ACR as the training loss. Instead, they optimize various surrogate objectives, and finally evaluate the model with ACR. Thus, it is unclear whether and to what extent the design of training algorithms is affected by the ACR metric. We now empirically quantify the effect, confirming that the theoretical weakness of ACR has introduced strong selection biases in existing RS training algorithms. Specifically, we show that SOTA training strategies reduce $p_{A}$ of hard samples and put more weight (measured by gradient norm) on easy samples compared to Gaussian training.
Figure 3 shows the empirical cumulative distribution of $p_{A}$ for models trained with SOTA algorithms and Gaussian training. Clearly, for various $\sigma$, SOTA algorithms have higher density than Gaussian training at $p_{A}$ close to zero and one. While they gain more ACR due to the improvement on easy samples, hard samples are consistently underrepresented in the final model compared to Gaussian training. As a result, Gaussian training has higher $\mathbb{P}(p_{A}\geq 0.5)$ (clean accuracy), and SOTA algorithms exceed Gaussian training only when $p_{A}$ passes a certain threshold, i.e., when the certified radius is relatively large. This is problematic in practice, indicating that ACR does not properly measure the model’s ability even with a fixed certification setting. For example, a face recognition model could have a high ACR but consistently refuse to learn some difficult faces, which is not acceptable in real-world applications.
To further quantify the relative weight between easy and hard samples indicated by each training algorithm, we measure the average gradient $l_{2}$ norm of easy and hard samples for models trained with different algorithms and $\sigma=0.5$, as a proxy for the sample weight. Intuitively, samples with larger gradients contribute more to training and thus are more important for the final model. As shown in Table 1, Gaussian training puts less weight on easy samples than hard samples, which is natural as easy samples have smaller loss values. However, SOTA algorithms put more weight on easy samples compared to Gaussian training, e.g., the relative weight between easy and hard samples is 4.28 for the best ACR baseline, CAT-RS (Jeong et al., 2023), while for Gaussian training it is 0.45. This confirms that SOTA algorithms indeed prioritize easy samples over hard samples, consistent with the theoretical analysis in Section 4.2.
@@
Table 1: The average gradient $l_{2}$ norm of easy ($p_{A}>0.5$) and hard ($p_{A}<0.5$) samples for models trained with different algorithms and $\sigma=0.5$, along with their relative magnitude (easy / hard). The corresponding ACR is provided for reference.
@@ 5 Replicating the Progress in ACR
In Section 4, we find that ACR strongly prefers easy samples, and this bias has been introduced to the design of RS training strategies. This makes one wonder the question of whether explicitly focusing on easy data during training can effectively replicate the increase in ACR gained by years spent developing RS training strategies. In this section, we propose three intuitive modifications to the simplest Gaussian training to achieve this goal.
@@ 5.1 Discard Hard Data During Training
Samples with low $p_{A}$ contribute little to ACR, especially those with $p_{A}<0.5$ which have no contribution at all (Section 4.2). Further, as shown in Figure 3, more than 20% of the data has $p_{A}<0.5$ after training converges. Therefore, we propose to discard hard samples directly during training, so that they explicitly have no effect on the final training convergence. Specifically, given a warm-up epoch $E_{t}$ and a confidence threshold $p_{t}$, we discard all data samples with $p_{A}<p_{t}$ at epoch $E_{t}$. We fix the number of steps taken by gradient descent and re-iterate on the distilled dataset when necessary to minimize the difference in training budget. After the discard, Gaussian training also ignores hard inputs, similar to SOTA algorithms.
@@ 5.2 Data Reweighting with Certified Radius
ACR relates non-linearly to $p_{A}$, and the growth of the certified radius is much faster for easy samples. We account for this by reweighting the data points based on their certified radius. Specifically, we use the approximate (normalized) certified radius as the weight of the probability for every data point being sampled. We formulate the weight $w$ of every data point $x$ as:
@@
where $C$ is the count of correctly classified noisy samples and $p_{\rm min}$ is the reference probability threshold. Note that the estimation of $\hat{p}_{A}$ aligns with the certification. We clip the weight at 1 to avoid zero weights. To minimize computational overhead, we evaluate $\hat{p}_{A}$ every 10 epoch with $N=16$ and $\alpha=0.1$ throughout the paper. We set $p_{\rm min}=0.75$ because this is the minimum probability that has positive certified radius under this setting. The sampling weight curve is visualized in Figure 4. After the reweighting, ACR has an approximately linear relationship to $p_{A}$ when $w>1$, and easy samples are sampled more frequently than hard samples so that Gaussian training also improves $p_{A}$ further for easy samples.
@@ 5.3 Adaptive Attack on the Sphere
(a) Before adaptive attack
(b) After adaptive attack
@@
$\delta^{*}=\operatorname*{arg\,min}_{f(x+\delta)\neq c,\|\delta\|_{2}=\|\delta_{0}\|_{2}}\|\delta-\delta_{0}\|_{2},$
where $\delta_{0}$ is a random Gaussian noise sample. Note that when $x+\delta_{0}$ makes the base classifier misclassify, we have $\delta^{*}=\delta_{0}$, thus hard inputs are not affected by the adaptive attack. In addition, we remark that we do not constrain $\delta^{*}$ to be in the neighborhood of $\delta_{0}$ which is adopted by CAT-RS (Jeong et al., 2023); instead, we only maintain the $l_{2}$ norm of the noise, thus allowing the attack to explore a much larger space. This is because for every $\delta^{*}$ such that $\|\delta^{*}\|_{2}=\|\delta_{0}\|_{2}$, the probability of sampling $\delta^{*}$ is the same as $\delta_{0}$. We formalize this fact in Theorem 2.
@@
Assume $\delta_{1},\delta_{2}\in{\mathbb{R}}^{d}$ and $\delta_{1}\neq\delta_{2}$. If $\|\delta_{1}\|_{2}=\|\delta_{2}\|_{2}$, then $\mathbb{P}_{{\mathcal{N}}(0,\sigma^{2}I_{d})}(\delta_{1})=\mathbb{P}_{{\mathcal{N}}(0,\sigma^{2}I_{d})}(\delta_{2})$ for every $\sigma>0$.
The proof of Theorem 2 is deferred to Section A.2. Figure 5 visualizes the gradient norm distributions for different $p_{A}$ before and after the adaptive attack. We observe that the adaptive attack balances the gradient norm of easy and hard samples. Before the attack, the gradient norm of easy samples is much smaller than that of hard samples, while after the attack, the gradient norm of easy samples is amplified without interfering the gradient norm of hard samples. Therefore, with the adaptive attack, Gaussian training obtains a similar gradient norm distribution to SOTA algorithms, and it can find effective noise samples more efficiently. Pseudocode of the adaptive attack is shown in Algorithm 1, and more detailed description is provided in Section C.2.
Algorithm 2 Overall Training Procedure
  for epoch $=1$ to $N_{\rm epoch}$ do
[…]
  else
[…]
  end if
[…]
  end for
  end if
  end for
[…]
@@ 5.4 Overall Training Procedure
We now describe how the above three modifications are combined. At the beginning, we train the model with the Gaussian training (Cohen et al., 2019), which samples $m$ noisy points from the isometric Gaussian distribution uniformly at random and uses the average loss of noisy inputs as the training loss. When we reach the pre-defined warm-up epoch $E_{t}$, all data points with $p_{A}<p_{t}$ are discarded, and the distilled dataset is used thereafter, as described in Section 5.1. After this, we apply dataset reweighting and the adaptive attack to training. Specifically, every 10 epoch after $E_{t}$ (including $E_{t}$), we evaluate the model with the procedure described in Section 5.2 and assign the resulting sampling weight to each sample in the train set. In addition, we use the adaptive attack described in Section 5.3 to generate the noisy samples for training. The pseudocode is shown in Algorithm 2.
6 Experimental Evaluation
@@
We now evaluate our method proposed in Section 5 extensively. Overall, our method always achieves better ACR than SOTA methods, which indicates that focusing on easy data can effectively improve ACR.
Baselines. We compare our method to the following methods: Gaussian (Cohen et al., 2019), SmoothAdv (Salman et al., 2019), MACER (Zhai et al., 2020), Consistency (Jeong & Shin, 2020), SmoothMix (Jeong et al., 2021), and CAT-RS (Jeong et al., 2023). We always use the trained models provided by the authors if they are available and otherwise reproduce the results with the same setting as the original paper. We set $m=4$ for Gaussian training and our method, since this is the standard setting for SOTA methods (Jeong et al., 2023).
Main Result. Table 2 shows the ACR of different methods on CIFAR-10. Training details, including hyperparameters, are provided in Section C.1. We further include the results of independent runs in Section E.1 to validate the significance of our results. Our method consistently outperforms all baselines on ACR, which confirms the effectiveness of our modification to Gaussian training in increasing ACR. Further, our method successfully increases the certified accuracy at large radius, as we explicitly focus on easy inputs which is implicitly taken by other SOTA methods. Figure 6 further visualizes certified accuracy at different radii, showing that at small radii, SOTA methods (including ours) are systematically worse than Gaussian training, while at large radii, these methods consistently outperform Gaussian training. The success of our simple and intuitive modification to Gaussian training suggests that existing approaches do not fully converge on strategies exploiting the weakness of ACR, and that the field should re-evaluate RS training strategies with better metrics instead of discarding them completely.
@@ 7 Beyond the ACR Metric
This work rigorously shows that the weakness of the main evaluation metric ACR has introduced strong biases into the development of randomized smoothing training strategies. Therefore, the field has to seek better alternative metrics to evaluate the robustness of models under RS. When the certification setting (budget $N$ and confidence $\alpha$) is fixed, we suggest using the best certified accuracy at various radii. Specifically, models certified with smaller $\sigma$ tend to achieve higher certified accuracy at smaller radii, whereas models certified with larger $\sigma$ usually have higher certified accuracy at larger radii. To extensively illustrate the power of the developed training strategies and avoid such tradeoff and dependence on $\sigma$, we suggest reporting the best certified accuracy across $\sigma$ at each radius. In other words, this represents the setting where one first fixes an interested radius, and then try to develop the best model with the highest certified radius at the pre-defined radius. Notably, several studies already adopt this strategy (Carlini et al., 2022; Xiao et al., 2022), avoiding the weakness of ACR unintentionally. This evaluation is also consistent to the practice in deterministic certified training (Gowal et al., 2018; Mirman et al., 2018; Shi et al., 2021; Müller et al., 2023; Mao et al., 2023, 2024a, 2024b; Palma et al., 2023; Balauca et al., 2024). Furthermore, the community should also consider randomized smoothing methods with $\sigma$ as a variable rather than a hyperparameter, thus effectively removing the dependence on $\sigma$. While there are some preliminary works in this direction (Nurlanov et al., 2023), they usually break the theoretical rigor, thus more research is needed in this direction.
When the certification setting varies, which is common in applying randomized smoothing to practical tasks, metrics are required to reflect the power of RS training strategies under different certification budgets. To this end, we suggest reporting the empirical distribution of $p_{A}$ for a large $N$, e.g., $N=10^{5}$ as commonly adopted. This approximates the true distribution of $p_{A}$ nicely when $N$ is large, and a higher $p_{A}$ usually leads to a larger certified radius consistently under different certification budgets. To this end, as long as one model has larger $P(p_{A}\geq p)$ for every $p\in[0.5,1]$ than another model, we can claim the former is better consistently in different certification settings. In addition, future research may also report the results under different certification budgets, which can provide insights on the capability of the training strategies under different certification settings as well. With these metrics, practitioners can better understand the capability of RS training strategies under the desired certification setting. A case study is included in App. B to further demonstrate this.

### v4

@@ 4.3 Selection Bias in Existing RS Training Algorithms
Table 2: The average model parameters gradient $l_{2}$ norm of easy ($p_{A}>0.5$) and hard ($p_{A}<0.5$) samples for models trained with different algorithms and $\sigma=0.5$, along with their relative magnitude (easy / hard). The corresponding ACR is provided for reference.
We have shown that ACR strongly prefers easy samples in Section 4.2. However, since ACR is not differentiable with respect to the model parameters because it is based on counting, RS training strategies usually do not directly apply ACR as the training loss. Instead, they optimize various surrogate objectives, and finally evaluate the model with ACR. Thus, it is unclear whether and to what extent the design of training algorithms is affected by the ACR metric. We now empirically quantify the effect, confirming that the theoretical weakness of ACR has introduced strong selection biases in existing RS training algorithms. Specifically, we show that SOTA training strategies reduce $p_{A}$ of hard samples and put more weight (measured by gradient magnitude) on easy samples when compared to Gaussian training.
Figure 3 shows the empirical cumulative distribution of $p_{A}$ for models trained with SOTA algorithms and Gaussian training on CIFAR-10. Clearly, for various $\sigma$, SOTA algorithms have higher density than Gaussian training at $p_{A}$ close to zero and one. While they gain more ACR due to the improvement on easy samples, hard samples are consistently underrepresented in the final model compared to Gaussian training. The same results hold on ImageNet (Figure 11 in Section E.2). As a result, Gaussian training has higher $\mathbb{P}(p_{A}\geq 0.5)$ (clean accuracy), and SOTA algorithms exceed Gaussian training only when $p_{A}$ passes a certain threshold, i.e., when the certified radius is relatively large. This is problematic in practice, indicating that ACR does not properly measure the model’s ability even with a fixed certification setting. For example, a face recognition model could have a high ACR but consistently refuse to learn some difficult faces, which is not acceptable in real-world applications.
To further quantify the relative weight between easy and hard samples indicated by each training algorithm, we measure the average gradient $l_{2}$ norm of easy and hard samples for model parameters trained with different algorithms and $\sigma=0.5$, as a proxy for the sample weight. Intuitively, samples with larger gradients on parameters contribute more to training and thus are more important to the final model. As shown in Table 2, Gaussian training puts less weight on easy samples than hard samples, which is natural as easy samples have smaller loss values. However, SOTA algorithms put more weight on easy samples compared to Gaussian training, e.g., the relative weight between easy and hard samples is 4.28 for the best ACR baseline, CAT-RS (Jeong et al., 2023), while for Gaussian training it is 0.45. This confirms that SOTA algorithms indeed prioritize easy samples over hard samples, consistent with the theoretical analysis in Section 4.2.
@@ 5 Replicating the Progress in ACR
In Section 4, we find that ACR strongly prefers easy samples, and this bias has been introduced to the design of RS training strategies. This questions that whether explicitly focusing on easy data during training can effectively replicate the increase in ACR gained by years spent on the development of RS training strategies. In this section, we propose three intuitive modifications on the Gaussian training, answering this question affirmatively.
@@ 5.1 Discard Hard Data During Training
Samples with low $p_{A}$ contribute little to ACR, especially those with $p_{A}<0.5$ which have no contribution at all (Section 4.2). Further, as shown in Figure 3, 20%-50% of data points has $p_{A}<0.5$ after training converges on CIFAR-10. Therefore, we propose to discard hard samples directly during training, so that they explicitly have diminished effect on the final training convergence. Specifically, given a warm-up epoch $E_{t}$ and a confidence threshold $p_{t}$, we discard all data samples with $p_{A}<p_{t}$ at epoch $E_{t}$. We fix the number of steps taken by gradient descent and re-iterate on the distilled dataset when necessary to minimize the difference in training budget. After the discard, Gaussian training also ignores hard inputs, similar to SOTA algorithms.
@@ 5.2 Data Reweighting with Certified Radius
ACR relates non-linearly to $p_{A}$, and the growth of the certified radius is much faster when improving easy samples. We account for this by reweighing the data points based on their certified radius. Specifically, we use the approximate (normalized) certified radius as the weight of the data point, realized by controlling the sampling frequency of each data point. The weight $w$ of every data point $x$ is defined as:
@@
where $C$ is the count of correctly classified noisy samples and $p_{\rm min}$ is the reference probability threshold. Note that the estimation of $\hat{p}_{A}$ aligns with the certification. We clip the weight at 1 to avoid zero weights. To minimize computational overhead, we evaluate $\hat{p}_{A}$ every 10 epoch with $N=16$ and $\alpha=0.1$ throughout the paper. We set $p_{\rm min}=0.75$ because this is the minimum probability that has positive certified radius under this setting. The sampling weight curve is visualized in Figure 4. After the reweight, ACR has an approximately linear connection to $p_{A}$ when $w>1$, and easy samples are sampled more frequently than hard samples so that Gaussian training also improves $p_{A}$ further for easy samples.
@@ 5.3 Adaptive Attack on the Sphere
(a) Before adaptive attack
(b) After adaptive attack
@@
$\bm{\delta}^{*}=\operatorname*{arg\,min}_{f(x+\bm{\delta})\neq c,\|\bm{\delta}\|_{2}=\|\bm{\delta}_{0}\|_{2}}\|\bm{\delta}-\bm{\delta}_{0}\|_{2},$
where $\bm{\delta}_{0}$ is a random Gaussian noise sample. Note that when $x+\bm{\delta}_{0}$ makes the base classifier misclassify, we have $\bm{\delta}^{*}=\bm{\delta}_{0}$, thus hard inputs are not affected by the adaptive attack. In addition, we remark that we do not constrain $\bm{\delta}^{*}$ to be in the neighborhood of $\bm{\delta}_{0}$ which is adopted by CAT-RS (Jeong et al., 2023); instead, we only maintain the $l_{2}$ norm of the noise, thus allowing the attack to explore a much larger space. This is because for every $\bm{\delta}^{*}$ such that $\|\bm{\delta}^{*}\|_{2}=\|\bm{\delta}_{0}\|_{2}$, the value of the probability density function at $\bm{\delta}^{*}$ is the same as $\bm{\delta}_{0}$. We formalize this fact in Theorem 2.
@@
Assume $\bm{\delta}_{1},\bm{\delta}_{2}\in{\mathbb{R}}^{d}$ and $\bm{\delta}_{1}\neq\bm{\delta}_{2}$. If $\|\bm{\delta}_{1}\|_{2}=\|\bm{\delta}_{2}\|_{2}$, then $p_{{\mathcal{N}}(0,\sigma^{2}I_{d})}(\bm{\delta}_{1})=p_{{\mathcal{N}}(0,\sigma^{2}I_{d})}(\bm{\delta}_{2})$ for every $\sigma>0$, where $p_{{\mathcal{N}}(0,\sigma^{2}I_{d})}$ is the probability density function of the isometric Gaussian distribution with mean 0 and covariance $\sigma^{2}I_{d}$.
The proof of Theorem 2 is deferred to Section A.2. Figure 5 visualizes the gradient norm distributions for different $p_{A}$, before and after the adaptive attack. We observe that the adaptive attack balances the gradient norm of easy and hard samples. Before the attack, the gradient norm of easy samples is much smaller than that of hard samples, while after the attack, the gradient norm of easy samples is amplified without interfering the gradient norm of hard samples. Therefore, with the adaptive attack, Gaussian training obtains a similar gradient norm distribution to SOTA algorithms, and it can find effective noise samples more efficiently. Pseudocode of the adaptive attack is shown in Algorithm 1, and more detailed description is provided in Section C.3.
@@ 5.4 Overall Training Procedure
We now describe how the above three modifications are combined. At the beginning, we train the model with the Gaussian training (Cohen et al., 2019), which samples $m$ noisy points from the isometric Gaussian distribution uniformly at random and uses the average loss of noisy inputs as the training loss. When we reach the pre-defined warm-up epoch $E_{t}$, all data points with $p_{A}<p_{t}$ are discarded, and the distilled dataset is used thereafter, as described in Section 5.1. After this, we apply dataset reweight and the adaptive attack to training. Specifically, every 10 epoch after $E_{t}$ (including $E_{t}$), we evaluate the model with the procedure described in Section 5.2 and assign the resulting sampling weight to each sample in the train set. In addition, we use the adaptive attack described in Section 5.3 to generate the noises used for training. The pseudocode is shown in Algorithm 2 in Section C.1.
@@
6 Experimental Evaluation
We now evaluate our method proposed in Section 5 extensively. Overall, simply focusing on easy samples is enough to replicate the advances in RS training strategies, sometimes even surpassing existing SOTA algorithms, e.g., , on CIFAR-10. Experiment details, including hyperparameters, are provided in Section C.1.
Baselines. We compare our method to the following methods: Gaussian (Cohen et al., 2019), SmoothAdv (Salman et al., 2019), MACER (Zhai et al., 2020), Consistency (Jeong & Shin, 2020), SmoothMix (Jeong et al., 2021), and CAT-RS (Jeong et al., 2023). We always use the trained models provided by the authors if they are available and otherwise reproduce the results with the same setting as the original paper. For CIFAR-10, we set $m=4$ for Gaussian training and our method, since this is the standard setting for SOTA methods (Jeong et al., 2023).
$\sigma$ Methods ACR 0.0 0.5 1.0 1.5 2.0 2.5 3.0 3.5 0.25 Gaussian 0.476 66.7 49.4 0.0 0.0 0.0 0.0 0.0 0.0 SmoothAdv 0.532 65.3 55.5 0.0 0.0 0.0 0.0 0.0 0.0 Ours 0.529 64.2 55.8 0.0 0.0 0.0 0.0 0.0 0.0 0.5 Gaussian 0.733 57.2 45.8 37.2 28.6 0.0 0.0 0.0 0.0 SmoothAdv 0.824 53.6 49.4 43.3 36.8 0.0 0.0 0.0 0.0 Consistency 0.822 55.0 50.2 44.0 34.8 0.0 0.0 0.0 0.0 SmoothMix 0.846 54.6 50.0 43.4 37.8 0.0 0.0 0.0 0.0 Ours 0.842 55.5 48.6 43.8 37.6 0.0 0.0 0.0 0.0 1.0 Gaussian 0.875 43.6 37.8 32.6 26.0 19.4 14.8 12.2 9.0 SmoothAdv 1.040 40.3 37.0 34.0 30.0 26.9 24.6 19.7 15.2 Consistency 0.982 41.6 37.4 32.6 28.0 24.2 21.0 17.4 14.2 SmoothMix 1.047 39.6 37.2 33.6 30.4 26.2 24.0 20.4 17.0 CAT-RS 1.071 43.6 38.2 35.2 30.8 26.8 24.0 20.2 17.0 Ours 1.042 41.0 37.4 33.6 31.0 27.0 23.2 19.4 15.8
Table 4: Comparison of certified test accuracy (%) at different radii and ACR on ImageNet. This table is structured in the same way as Table 3.
Main Result. Table 3 shows the ACR of different methods on CIFAR-10. We further include the results of independent runs in Section E.1 to validate the significance of our results. Our method consistently outperforms all baselines on ACR, which confirms the effectiveness of our modification to Gaussian training in increasing ACR. Further, our method successfully increases the certified accuracy at large radius, as we explicitly focus on easy inputs which is implicitly taken by other SOTA methods. Figure 6 further visualizes certified accuracy at different radii, showing that at small radii, SOTA methods (including ours) are systematically worse than Gaussian training, while at large radii, these methods consistently outperform Gaussian training.
Table 4 shows the results on ImageNet, which is organized in the same way as Table 3. Our method replicates a large portion of the ACR advances. Concretely, for $\sigma=0.25$, the proposed method recovers (0.529 - 0.476) / (0.532 - 0.476) = 94.6% of the advance; for $\sigma=0.5$, it recovers (0.842 - 0.733) / (0.846 - 0.733) = 96.5%; for $\sigma=1.0$, it recovers (1.042 - 0.875) / (1.071 - 0.875) = 85.2%. This confirms that our results generalize across datasets.
The success of our simple and intuitive modification to Gaussian training suggests that existing approaches do not fully converge on strategies exploiting the weakness of ACR, and that the field should re-evaluate RS training strategies with better metrics instead of discarding them completely.
@@ 7 Beyond the ACR Metric
This work rigorously shows that the weakness of the main evaluation metric ACR has introduced strong biases into the development of RS training strategies. Therefore, the field has to seek better alternative metrics to evaluate the robustness of models under RS. When the certification setting (budget $N$ and confidence $\alpha$) is fixed, we suggest using the best certified accuracy at various radii. Specifically, models certified with smaller $\sigma$ tend to achieve higher certified accuracy at smaller radii, whereas models certified with larger $\sigma$ usually have higher certified accuracy at larger radii. To extensively illustrate the power of the developed training strategies and avoid such tradeoff and dependence on $\sigma$, we suggest reporting the best certified accuracy across $\sigma$ at each radius. In other words, this represents the setting where one first fixes an interested radius, and then try to develop the best model with the highest certified accuracy at the pre-defined radius. Notably, several studies already adopt this strategy (Carlini et al., 2023; Xiao et al., 2022), avoiding the weakness of ACR unintentionally. This evaluation is also consistent to the practice in deterministic certified training (Gowal et al., 2018; Mirman et al., 2018; Shi et al., 2021; Müller et al., 2023; Mao et al., 2023, 2024a, 2024b; De Palma et al., ; Balauca et al., 2024). Furthermore, the community should also consider RS methods with $\sigma$ as a variable rather than a hyperparameter, thus effectively removing the dependence on $\sigma$. While there are some preliminary works in this direction (Alfarra et al., 2022; Wang et al., 2021), they usually break the theoretical rigor (Súkeník et al., 2022). Rencent works propose sound methods to certify with various $\sigma$ (Súkeník et al., 2022; Jeong & Shin, 2024). However, room for improvement remains, and further research is needed.
When the certification setting varies, which is common in applying RS to practice, metrics are required to reflect the power of RS training strategies under different certification budgets. To this end, we suggest reporting the empirical distribution of $p_{A}$ for a large $N$, e.g., $N=10^{5}$ as commonly adopted. This approximates the true distribution of $p_{A}$ nicely when $N$ is large, and a higher $p_{A}$ usually leads to a larger certified radius consistently under different certification budgets. To this end, as long as one model has larger $P(p_{A}\geq p)$ for every $p\in[0.5,1]$ than another model, we can claim the former is better consistently in different certification settings. A case study is included in App. B to further demonstrate the usage and properties of the empirical distribution of $p_{A}$.